OpenStudy (anonymous):

In Pensacola in June, high tide was at noon. The water level at high tide was 12 feet and 2 feet at low tide. Assuming the next high tide is exactly 12 hours later and that the height of the water can be modeled by a cosine curve, find an equation for water level in June for Pensacola as a function of time (t).

OpenStudy (anonymous):

f(t) = 12 cospi over 2t + 5 f(t) = 5 cospi over 2t + 12 f(t) = 5 cospi over 6t + 7 f(t) = 7 cospi over 6t + 12

OpenStudy (anonymous):

@cp9454

ganeshie8 (ganeshie8):

high tide = 12 feet low tide = 2 feet so, peak to peak = 12-2 = 10 amplitude = 10/2 = 5

ganeshie8 (ganeshie8):

In light of above information, which two options can you eliminate ?

OpenStudy (anonymous):

C and..... @ganeshie8

OpenStudy (anonymous):

B?

ganeshie8 (ganeshie8):

do you mean, the answer is betwee B and C ?

OpenStudy (anonymous):

I guess lol

ganeshie8 (ganeshie8):

then you're correct :) it is between B and C

ganeshie8 (ganeshie8):

`the next high tide is exactly 12 hours later ` that means, period = 12

ganeshie8 (ganeshie8):

whats the period of cosine function in option B ?

OpenStudy (anonymous):

12?

ganeshie8 (ganeshie8):

how ?

OpenStudy (anonymous):

I guessed, cause the last number is 12 in their equation

ganeshie8 (ganeshie8):

haha thats a very good guess, but the last number does not represent period - it represents the vertical shift

OpenStudy (anonymous):

so then how would I figure out what the period of cosine function is?

ganeshie8 (ganeshie8):

\[\large y = A\cos(Bt ) + C\] period = \(\large \dfrac{2\pi}{B}\)

OpenStudy (anonymous):

Answer is C

ganeshie8 (ganeshie8):

Yep !